Consider the following IS–LM model:


C = 400 + 0.25YD


I = 300 + 0.25Y − 1500i


G = 600


T = 400


(M/P)d = 2Y − 12 000i


M/P = 3000


a. Derive the IS relation. (Hint: You want an equation with Y on the left side and everything else on the right.)


b. Derive the LM relation. (Hint: It will be convenient for later use to rewrite this equation with i on the left side and everything else on the right.)


1
Expert's answer
2021-11-15T17:51:43-0500

a.

In goods market, Y = C + I + G

Y = 400 + 0.25(Y - 400) + 300 + 0.25Y - 1500i + 600 [as Yd = Y - T]

Y = 1300 + 0.25Y - 100 + 0.25Y - 1500i

0.5Y = 1200 - 1500i

Y = 2400 - 3000i..........(IS curve)

b.

In money market, (M/P)d = (M/P)s

2Y - 12000i = 3000

2Y = 3000 + 12000i

i = (2Y - 3000) / 12000

i = (Y - 1500) / 6000.........(LM curve)